# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ III-A Applications of Wireless Positioning
Wireless positioning has emerged as a transformative technology, enabling a wide range of applications that improve efficiency, safety, and user experience across various domains. From intelligent transportation systems to healthcare and beyond, the ability to accurately determine the location of devices and objects has unlocked significant advancements in modern technologies. In this subsection, we will introduce the importance of wireless positioning from several representative application scenarios.
III-A1 Intelligent Transportation
@@ III-A2 Autonomous Driving
III-A3 Extended Reality
In XR applications, precise positioning of devices and users is crucial for creating seamless and immersive experiences. Whether in augmented reality (AR) navigation [63], virtual reality (VR) training simulations [64, 65, 66], or multiplayer gaming environments [67], wireless positioning ensures accurate tracking of movements and spatial relationships. This enhances the realism of interactive simulations and allows for seamless transitions between indoor and outdoor environments in XR-based applications.
III-A4 Indoor Tracking and Navigation
@@ III-A7 Security and Surveillance
III-A8 Sports and Motion Analysis
@@ III-A9 Healthcare
III-B Emerging Wireless Technologies
Emerging wireless technologies are reshaping the landscape of wireless positioning by introducing novel architectures and signal propagation characteristics that enhance accuracy, scalability, and reliability. Key advancements include cell-free networks, massive multiple-input multiple-output (MIMO), non-terrestrial networks (NTNs), THz communications, and RIS. Subsequently, we will introduce several key technologies.
III-B1 Massive MIMO
Massive MIMO technology, with its deployment of large-scale antenna arrays at base stations (BSs), significantly improves spatial resolution and beamforming precision. This enhances the ability to estimate parameters such as the angles-of-arrival (AOA), direction-of-arrival (DOA), angle-of-departure (AOD), and TOA. The high spatial diversity provided by massive MIMO also mitigates the effects of multipath propagation, making it a cornerstone for precise positioning in 5G and beyond. Additionally, Flexible Location MIMO [83], based on fluid antennas and movable antennas, represents a novel antenna paradigm. These antennas can move dynamically to avoid interference and obstacles, ensuring fast and reliable wireless transmission.
III-B2 Cell-Free
Cell-free networks utilize a distributed architecture where a dense network of access points (APs) collaborates to serve users [84]. This architecture mitigates inter-cell interference and provides seamless coverage, enhancing positioning accuracy, especially in dense urban or indoor environments. By exploiting the joint processing of signals from multiple APs, cell-free networks enable highly precise multi-AP positioning [84]. Additionally, the system’s inherent flexibility supports adaptive coverage and robust performance in dynamic scenarios.
III-B3 Non-Terrestrial Networks
NTNs, encompassing satellite constellations, high-altitude platforms, and UAVs, extend wireless connectivity to remote and underserved areas [85]. These networks complement terrestrial systems by providing wide-area coverage and redundancy. For positioning, NTNs offer unique advantages, such as a clear LOS and large coverage areas, enabling applications in maritime, aviation, and disaster recovery scenarios.
III-B4 Terahertz Networks
THz communication operates in the spectrum above 100 GHz, offering ultra-wide bandwidths that allow high-resolution parameter estimation. The short wavelength of THz signals enables precise localization through fine-grained TOA and AOA measurements. These networks are particularly suited for ultra-dense environments and applications requiring sub-centimeter positioning accuracy, such as industrial automation and XR [18].
III-B5 reconfigurable intelligent surfaces
RIS technology uses programmable metasurfaces to manipulate electromagnetic waves, creating controllable propagation environments. By dynamically adjusting reflection, refraction, or scattering, RIS can enhance signal quality, and improve positioning accuracy in challenging environments. For example, RIS can redirect signals to establish virtual LOS paths or amplify weak signals in dense urban areas [86]. The integration of RIS with AI-driven positioning systems promises unprecedented levels of control and precision, though challenges such as hardware scalability and real-time optimization remain.
III-B6 Near-field communication
Near-field Communication enhances wireless positioning by leveraging spherical wavefronts in close proximity environments, enabling precise parameter estimation like phase and amplitude variations [87]. This improves accuracy in technologies such as massive MIMO and RIS, making it ideal for applications like indoor navigation and industrial automation. However, near-field communication increases computational complexity and requires advanced models to capture wavefront characteristics, presenting both opportunities and challenges for next-generation wireless networks.
The integration of these technologies significantly enhances wireless positioning capabilities. By leveraging their unique characteristics, such as improved spatial resolution in massive MIMO, flexibility in RIS, and wide coverage in NTNs, these advancements address the limitations of conventional systems. Furthermore, combining these technologies with AI holds promise for achieving robust, scalable, and ultra-precise positioning in next-generation wireless networks.
III-C Basics of Wireless Positioning
Wireless positioning technology fundamentally involves the estimation of a device or object’s location through the measurement of the propagation characteristics of wireless signals between a transmitter and a receiver. Path loss is a key and intuitive factor in this process, which describes the attenuation of signal power over distance [88]. Path loss can be expressed using models like the free-space path loss equation:
@@ III-C Basics of Wireless Positioning
where $PL(d)$ represents the path loss as a function of distance $d$, $PL(d_{0})$ is the path loss at a reference distance $d_{0}$, and $n$ is the path loss exponent, which depends on the environment.
Based on path loss, signal strength inherently reflects the spatial relationship between the transmitter and receiver. This characteristic makes received signal strength (RSS) metrics, such as received signal strength indicator (RSSI) and reference signal received power (RSRP), valuable measurements for wireless positioning. RSSI represents a measure of the total power present in a received radio signal, encompassing both the desired signal and any interference or noise [89]. It is widely used in wireless technologies such as WiFi, Zigbee, and Bluetooth. On the other hand, RSRP is a metric specific to Long-term Evolution (LTE) and 5G networks [90]. It measures the power of the reference signal transmitted by BSs, providing a more precise and isolated indicator of signal quality and strength for cellular systems. These metrics allow wireless systems to infer distances between devices and BSs, forming the foundation for various positioning techniques. By relating measured RSSI or RSRP to known transmission power levels via path loss models, the distance between the transmitter and receiver can be estimated. These distance estimates form the foundation of range-based positioning techniques, such as trilateration [91], which will be discussed later.
However, in the practical system, the use of RSSI and RSRP is affected by factors such as multipath propagation, environmental obstacles and noise, which may introduce errors into the distance estimation [92]. To better characterize the wireless transmission environment, a more detailed metric, channel state information (CSI), is often employed [93]. CSI provides a comprehensive characterization of the wireless channel response between the transmitter and receiver, including both amplitude and phase information for each subcarrier in the frequency domain. Unlike RSSI and RSRP, which are scalar measures, CSI is a multidimensional metric that captures the channel’s response across frequencies or subcarriers. Additionally, the detailed frequency-domain information provided by CSI allows for advanced modeling techniques, such as ML-based approaches, to further enhance positioning accuracy in dynamic and challenging environments [94].
@@
As an illustrative example, we consider a MIMO-orthogonal frequency-division multiplexing (OFDM) system to model CSI in a wireless positioning system. As shown in Fig. 5, in the MIMO-OFDM system, the BS is equipped with a uniform plane array, comprising $N_{a}$ antennas with $N_{a}^{r}$ in each row and $N_{a}^{c}$ in each column. The user equipment (UE) is equipped with a single omnidirectional antenna. The system contains $N_{s}$ subcarriers, each with a bandwidth of $B_{s}$. The locations of the BS and the UE can be represented as $\bm{p}^{ue}$ and $\bm{p}^{bs}$ respectively. The channel frequency response between the BS and the UE for subcarrier $s$ can be written as [95]
@@
where $f_{s}$ is the subcarrier frequency, $\alpha_{l}$ and $\tau_{l}$ denote the complex gain and delay of path $l$, and $\mathbf{a}(\theta_{l},\varphi_{l})$ is the two-dimensional array response matrix. For the LOS path, $\tau_{1}$ and $(\theta_{1},\varphi_{1})$ represent the delay and AOA, respectively.
Specifically, $\mathbf{a}(\theta_{l},\varphi_{l})$ can be expressed as
@@
where $\bm{p}^{ue}$ and $\hat{\bm{p}}^{ue}$ represent the estimated position and the true position, respectively. However, this is a challenging problem due to the complexity of the wireless propagation environment, which includes factors such as multipath propagation, noise, and NLOS conditions.
To find an optimal $\mathcal{F}(\cdot)$, existing research generally adopts two main approaches. The first approach relies on the geometric relationships between the transmitter and the device, focusing on extracting geometric features like TOA and AOA from the CSI. These features are then utilized to calculate the UE’s position based on geometric principles.
The second approach uses artificial intelligence algorithms for direct end-to-end localization without explicitly estimating intermediate geometric parameters. For instance, in fingerprint-based positioning, raw CSI data or its derived features are used as inputs to AI models, which then output the position. These data-driven methods learn complex mapping functions between CSI and location, enabling high-precision positioning without relying on geometric computations.
@@
III-C1 Ranging-based Positioning
@@
High-precision positioning can be achieved by estimating the distances between a device and multiple BSs. With geometric relationships, accurate positioning is possible when distance estimations to at least three BSs are available. As shown in Fig. 6(a), the intersection of circles representing these distances identifies the device’s location. Adding more BSs and improving distance estimation accuracy further enhance positioning precision.
As previously discussed, one way to measure the distance between the BSs and the device is by fitting the relationship between the RSS and the distance, thus allowing distance estimation using RSS [96, 97, 98, 99, 100]. The same principles apply to methods based on RSRP in cellular systems and RSSI in WiFi systems. However, the propagation characteristics of wireless signals vary across different frequency bands, and the wireless propagation environment fluctuates over time. These factors make it challenging to determine an accurate path loss function. Factors such as multipath propagation, interference, and NLOS conditions further degrade the accuracy of RSS-based methods. Consequently, although RSS-based positioning is simple and cost-effective, it is not easy to achieve high-precision positioning.
To overcome the limitations of RSS-based methods, TOA estimation provides a more precise approach. TOA, also referred to as time-of-flight (TOF) [14], is used to measure the time taken for a wireless signal to propagate from the transmitter to the receiver. Based on the TOA measurement, the distance $d$ can then be calculated as $d=c\cdot t_{toa}$, where $t_{toa}$ is the signal propagation time, equal to the propagation delay of the LOS path $\tau_{1}$, and $c\approx 3\times 10^{8}$ m/s is the speed of light.
In practice, to estimate the TOA, BS transmits a predefined pilot signal, which is received by the device. The device performs channel estimation using the received signal and then estimates the TOA from the estimated channel [101, 102, 103, 104]. The TOA estimation process can be expressed as a function $\mathcal{F}_{toa}(\cdot)$ that maps the estimated channel $\hat{\mathbf{H}}$ to the TOA:
$t_{toa}=\mathcal{F}_{toa}(\hat{\mathbf{H}}).$ (8)
$\hat{\mathbf{H}}$ represents the estimated CSI, and $\mathcal{F}_{toa}(\cdot)$ denotes signal processing algorithms used to extract the TOA from the channel response. The accuracy of TOA-based distance estimation is influenced by several factors, including the pilot signal bandwidth, signal-to-noise ratio (SNR), and the sampling rate of the receiver. The Cramer-Rao lower bound (CRLB) can be used to describe the variance of the TOA estimation error as a function of the bandwidth $B$ and SNR [104, 105, 106, 107]:
$\text{Var}(t_{toa})\propto\frac{1}{B^{2}\cdot\text{SNR}}.$ (9)
This indicates that increasing the bandwidth and SNR enhances the accuracy of TOA estimation, albeit at the cost of requiring more advanced hardware and signal processing.
To improve TOA estimation accuracy, advanced algorithms like multiple signal classification (MUSIC) [108] and estimation of signal parameters via rotational invariance techniques (ESPRIT) [109] are widely used. MUSIC identifies peaks in the signal’s spatial spectrum, while ESPRIT leverages the rotational invariance of the signal subspace for precise estimation. However, these algorithms still face the challenge of robustness in different environments. Additionally, TOA-based positioning requires precise clock synchronization between the BS and UEs [110, 111, 112, 113], as clock misalignment introduces an unknown offset in TOA measurements. This offset introduces a systematic error in the distance estimation. Therefore, the measured TOA can be expressed as
$t_{toa}^{\text{measured}}=t_{toa}+\Delta t_{clk}+e_{toa},$ (10)
where $\Delta t_{clk}$ and $e_{toa}$ represents the clock synchronization error and estimation error.
To mitigate the effects of clock synchronization errors, alternative methods such as round-trip time (RTT) [114, 115, 116] estimation are often employed. RTT measures the total round-trip delay of the signal, effectively eliminating the need for synchronization between the BS and the device. Since the RTT is calculated using the clock in the transmitter, it effectively cancels out the impact of clock misalignment.
Moreover, when BSs are synchronized with each other, TDOA-based positioning technique [117, 118, 119] provides an effective alternative to address the lack of synchronization between the BS and the device. TDOA measures the difference in arrival times of a signal at multiple BSs. For example, the TDOA can be expressed as
$\Delta t_{tdoa}=$ $\ t_{toa,i}-t_{toa,j}$
$=$ $\ t_{toa,i}+\Delta t_{clk,i}+e_{toa,i}-t_{toa,j}+\Delta t_{clk,j}+e_{toa,j}$
$=$ $\ t_{toa,i}-t_{toa,j}+e_{tdoa}^{*},$ (11)
where $t_{toa,i}$ and $t_{toa,j}$ are the TOAs from the device to BS $i$ and BS $j$, respectively, and $e_{tdoa}^{*}$ represents the combined error. In this case, the error caused by the desynchronization can be eliminated. Similar to the TOA-baed positioning, TDOA-based positioning, as shown in Fig. 6(b), can draw a hyperbola based on the TDOA between each pair of BSs and the device. When the number of BSs exceeds three, the region where the device is located can be determined. Similar to TOA-based positioning techniques, the more BSs involved in the positioning process and the more accurate the TDOA estimation, the higher the positioning accuracy of the device. In cellular positioning systems (LTE and 5G), TDOA-based techniques are also referred to as observed time difference of arrival (OTDOA) [104, 105, 120].
III-C2 Angle-based positioning
@@
Angle-based positioning techniques estimate the location of a device by determining angular parameters, specifically the AOA [121, 122, 123] and AOD [124, 125, 126], using the spatial characteristics of signals received or transmitted by antenna arrays.
The AOA indicates the direction from which a signal is received at the receiver, typically measured using uplink signals, while the AOD denotes the direction in which a signal is transmitted from the source, generally derived from downlink channels. For AOA estimation, For 3D positioning using AOA, it is crucial to estimate the 3D angular parameters, including the elevation angle $\theta_{1}$ and azimuth angle $\phi_{1}$ of the LOS path. Based on the channel model provided above, the AOA measurement error can be expressed as
$[\theta_{1},\phi_{1}]=\mathcal{F}_{aoa}(\hat{\mathbf{H}}).$ (12)
@@
For 2D positioning, we only need to estimate $\theta_{1}$. The CRLB for the 2D angle estimates can be expressed as [107]
@@
Similar to ranging-based positioning techniques, as illustrated in Fig. 6(c), angle-based positioning estimates the position of a device by using the spatial direction between the device and multiple BSs. A straightforward approach involves beam scanning, where the beam direction with the maximum received power is selected as the AOA estimation result. In addition, advanced algorithms, such as MUSIC [127, 128] and ESPRIT [129, 130], are widely employed. These methods leverage the spatial characteristics of the received signal, enabling angular resolution that exceeds the physical limitations of the antenna array.
The primary advantage of angle-based positioning is its independence from time synchronization between the BSs and the device, which is a crucial requirement for time-based methods like TOA and TDOA. However, the performance of angle-based positioning greatly depends on the angular resolution and precision of the antenna arrays, as well as the algorithms used for AOA and AOD estimation. Higher resolution arrays and advanced signal processing techniques can significantly improve positioning accuracy but also demand more advanced hardware and increased computational resources.
III-C3 Fingerprint-based Positioning
@@
Fingerprint-based positioning establishes a mapping between wireless network measurements and spatial locations by constructing an offline database or training AI models. It involves deriving a mapping function $\mathcal{F}{fp}(\cdot)$ from an offline dataset $D{train}={\bm{x}_{i},\bm{p}_{i}}$, where $\bm{x}_{i}$ represents wireless fingerprint features and $\bm{p}i$ denotes the corresponding UE position. The function $\mathcal{F}{fp}(\cdot)$ is obtained through database creation or AI model training.
As depicted in Fig. 6(d), fingerprint-based positioning typically consists of two main phases: the offline phase and the online phase. During the offline phase, a detailed survey of the environment is performed to collect wireless signal parameters, which are then used to construct a wireless fingerprint database, also known as a radio map. The key steps involved are as follows:
  • Offline Data Collection: Wireless signal characteristics, such as RSSI [131, 132, 133], RSRP [134, 135, 136, 137, 138, 139], CSI [140, 141, 142, 143, 144], or TOA measurements [145, 146], are collected at predefined reference points (RPs) within the target area. These RPs are typically arranged in a grid or sampled randomly.
  • Fingerprint Feature Extraction: The collected raw data is processed to extract meaningful features, such as average signal strength or variance. Alternatively, the raw signal data itself can also be used directly as a fingerprint.
  • Fingerprint Database Creation: The extracted features for each RP are compiled into a fingerprint database, denoted as $D_{train}$. Each database entry comprises a pair, $\bm{x}_{i}$ (signal features) and $\bm{p}i$ (corresponding RP coordinates), forming a virtual map of the fingerprint. Beyond traditional database construction [147], AI/ML models can be utilized to learn a mapping function $\mathcal{F}{fp}(\cdot)$ that directly correlates input features to output positions, serving as an alternative to conventional databases.
In the online phase, the goal is to determine the position of an unknown device by comparing its real-time signal measurements with the pre-established fingerprint database. Therefore, For a UE with fingerprint features $\bm{x}$, its estimated position can be expressed as
$\hat{\bm{p}}=\mathcal{F}_{pf}(\bm{x}).$ (14)
This position estimation process includes:
  • Real-Time Signal Measurement: The device requiring positioning continually collects wireless data at its current position, extracting the wireless fingerprint features.
  • Feature Matching and Position Estimation: Various matching algorithms are used to compare current signal measurements against entries in the fingerprint database. Location is estimated by identifying the RP whose fingerprints best match the observed signals. If an AI/ML model serves as the fingerprint mapping, the process involves inputting the extracted wireless features into the AI/ML model to output the estimated position directly.
  • Position Result Output: The result of the matching or model output is a predicted position, typically expressed as coordinates or descriptive position labels.
@@
III-C4 Channel Charting based Positioning
Channel charting is a novel approach leveraging CSI to enable the pseudo-positioning of users by modeling relationships between different channels [148]. It learns a mapping from CSI to a lower-dimensional channel chart, where distances in the latent space represent dissimilarity metrics between channels [149, 150, 151, 152]. These distances reflect the proximity of devices in the actual propagation environment, with higher channel similarity corresponding to closer proximity on the channel chart [148]. Besides positioning, channel charting based techniques have been explored in various applications, such as SNR prediction [153], pilot assignment [154, 155], pilot reuse [156], beam tracking [157] and wireless resource optimization [158].
In contrast to fingerprint-based methods, which depend on pre-collected wireless fingerprints and external RPsto achieve absolute positioning, channel charting employs self-supervised learning to reduce the dimensionality of CSI data. Its objective is to uncover latent spatial relationships while minimizing the need for labeled data.
Channel charting, a form of manifold learning, encompasses classical methods such as multidimensional scaling (MDS), Sammon mapping [159], and t-SNE [160], alongside deep learning approaches like Siamese Neural Networks and triplet-based dimensionality reduction.
In Siamese-based channel charting systems, dissimilarity metrics define relationships between CSI samples in the latent space. The channel charting learning process aims to ensure that latent space distances preserve the high-dimensional relationships captured by the dissimilarity metrics. A commonly used MDS-based loss function can be given by [150]

### v2

@@ II-A Applications of Wireless Positioning
Wireless positioning has emerged as a transformative technology, enabling a wide range of applications that improve efficiency, safety, and user experience across various domains. With the integration of AI and emerging wireless technologies such as RIS [24], Massive multiple-input multiple-output (MIMO) [25], THz communications [17], non-terrestrial networks (NTN) [26], near-field communications [27], and cell-free networks [28], seamless and high-precision wireless positioning is poised to drive significant advancements across various industries. In this subsection, we introduce the importance of wireless positioning through several representative application scenarios.
II-A1 Intelligent Transportation
@@ II-A2 Autonomous Driving
II-A3 Extended Reality
In XR applications, precise positioning of devices and users is crucial for creating seamless and immersive experiences. Whether in augmented reality (AR) navigation [36], virtual reality (VR) training simulations [37], or multiplayer gaming environments [38], wireless positioning ensures accurate tracking of movements and spatial relationships. This enhances the realism of interactive simulations and allows for seamless transitions between indoor and outdoor environments in XR-based applications.
II-A4 Indoor Tracking and Navigation
@@ II-A7 Security and Surveillance
II-A8 Sports and Motion Analysis
@@ II-A9 Healthcare
II-B Basics of Wireless Positioning
Wireless positioning estimates the location of a device by analyzing the propagation characteristics of wireless signals between transmitters and receivers. A fundamental concept is path loss, which describes the attenuation of signal power over distance [54] and can be modeled as:
@@ II-B Basics of Wireless Positioning
where $PL(d)$ represents the path loss as a function of distance $d$, $PL(d_{0})$ is the path loss at a reference distance $d_{0}$, and $n$ is the path loss exponent, which depends on the environment. Based on path loss, received signal strength (RSS), such as received signal strength indicator (RSSI), and reference signal received power (RSRP), reflects this spatial relationship and thus forms the basis of range-based positioning. RSSI measures total received signal power, including noise and interference [55], and is common in WiFi, Zigbee, and Bluetooth systems. RSRP, used in Long-Term Evolution (LTE)/5G [56], isolates the power of the reference signal, offering better precision for cellular positioning. Using known transmission power and path loss models, distances can be inferred and fed into techniques like trilateration [57]. However, RSSI/RSRP are sensitive to multipath, shadowing, and noise [58], limiting their accuracy. To better model the wireless channel, channel state information (CSI) is used [59]. Unlike scalar RSSI/RSRP, CSI provides multidimensional amplitude and phase information for each subcarrier, enabling advanced signal modeling, especially for AI-driven positioning [60].
@@
As an illustrative example, we consider a MIMO-orthogonal frequency-division multiplexing (OFDM) system to model CSI in a wireless positioning system. As shown in Fig. 3, the BS is equipped with a uniform plane array, comprising $N_{a}$ antennas with $N_{a}^{r}$ in each row and $N_{a}^{c}$ in each column. The user equipment (UE) is equipped with a single omnidirectional antenna. The system contains $N_{s}$ subcarriers, each with a bandwidth of $B_{s}$. The locations of the BS and the UE can be represented as $\bm{p}^{ue}$ and $\bm{p}^{bs}$ respectively. The channel frequency response between the BS and the UE for subcarrier $s$ can be written as [61]
@@
where $f_{s}$ is the subcarrier frequency, $\alpha_{l}$ and $\tau_{l}$ denote the complex gain and delay of path $l$, and $\mathbf{a}(\theta_{l},\varphi_{l})$ is the two-dimensional array response matrix. For the LOS path, $\tau_{1}$ and $(\theta_{1},\varphi_{1})$ represent the delay and angles-of-arrival (AOA), respectively. Specifically, $\mathbf{a}(\theta_{l},\varphi_{l})$ can be expressed as
@@
where $\bm{p}^{ue}$ and $\hat{\bm{p}}^{ue}$ represent the true position and the estimated position, respectively. However, this is a challenging problem due to the complexity of the wireless propagation environment, which includes factors such as multipath propagation, noise, and NLOS conditions. Two main approaches exist:
  • Model-based: Model-based solutions extract geometric features (e.g., TOA, AOA) from CSI, then estimate the location via geometric relationships.
  • Data-driven (AI-based): AI-based solutions use CSI directly as input to train models (e.g., deep learning) that learn the CSI-to-location mapping end-to-end.
@@
II-B1 Ranging-based Positioning
@@
High-precision positioning can be achieved by estimating the distances between a device and multiple BSs. With geometric relationships, accurate positioning is possible when distance estimations to at least three BSs are available. As shown in Fig. 4(a), the intersection of circles representing these distances identifies the device’s location. Adding more BSs and improving distance estimation accuracy further enhances positioning precision.
Among distance estimation methods, as previously mentioned, RSS-based techniques leverage the relationship between signal strength and distance [62, 63, 64, 65, 66]. However, their accuracy is limited by multipath propagation, interference, and environmental variability, making them unsuitable for high-precision applications. To address these limitations, TOA (also referred to as time-of-flight (TOF) [21]) estimation offers a more accurate alternative [21]. TOA measures the propagation delay of the LOS path, and the distance is calculated as $d=c\cdot t_{toa}$, where $t_{toa}$ is the signal propagation time, equal to the propagation delay of the LOS path $\tau_{1}$, and $c\approx 3\times 10^{8}$ m/s is the speed of light. TOA can be estimated from CSI using signal processing algorithms [67, 68, 69, 70], expressed as:
[…]
where $\mathcal{F}_{\text{toa}}(\cdot)$ is the estimation function and $\hat{\mathbf{H}}$ is the estimated channel. The estimation accuracy depends on signal bandwidth, signal-to-noise ratio (SNR), and sampling rate. According to the Cramer-Rao lower bound, the variance of the estimation error is inversely proportional to $B^{2}\cdot\text{SNR}$} [70, 71]. Advanced techniques, like multiple signal classification (MUSIC) [72] and estimation of signal parameters via rotational invariance techniques (ESPRIT) [73], can improve TOA accuracy, but their performance may degrade under multipath and dynamic environments. Moreover, TOA-based systems require tight clock synchronization between BS and UE, as clock offsets introduce systematic errors [74, 75, 76, 77] $t_{toa}^{\text{measured}}=t_{toa}+\Delta t_{clk}+e_{toa}$, where $\Delta t_{clk}$ and $e_{toa}$ represents the clock synchronization error and estimation error. To mitigate synchronization issues, round-trip time (RTT) estimation is often used, which eliminates the need for UE-BS synchronization by measuring total round-trip delay [78, 79, 80]. Since the RTT is calculated using the clock in the transmitter, it effectively cancels out the impact of clock misalignment.
Alternatively, when BSs are synchronized, TDOA offers another solution [81, 82, 83]. By subtracting the TOA measurements from two BSs, TDOA inherently eliminates clock offset errors, thus relaxing the need for UE-side synchronization. Based on TDOA, the algorithms perform positioning via hyperbolic intersections, as shown in Fig. 4(b). This method, also known as observed TDOA (OTDOA) in LTE/5G systems [84], avoids the need for UE-side synchronization and benefits from increasing BS count and estimation accuracy.
II-B2 Angle-based positioning
@@
Angle-based positioning techniques estimate the location of a device by determining angular parameters, specifically the AOA [85, 86, 87] and angle-of-departure (AOD) [88, 89, 90], using the spatial characteristics of signals received or transmitted by antenna arrays. AOA represents the incoming direction of a signal, typically estimated from uplink signals, while AOD refers to the signal’s transmission direction, often derived from downlink. For 3D positioning, both elevation angle $\theta_{1}$ and azimuth angle $\phi_{1}$ of the LOS path are required, and can be extracted from CSI as:
[…]
@@
For 2D positioning, we only need to estimate $\theta_{1}$. The estimation accuracy is bounded by the Cramer-Rao lower bound (CRLB) [91]:
@@
Similar to ranging-based methods, angle-based positioning infers location by intersecting spatial direction estimates from multiple BSs, as shown in Fig. 4(c). A simple method involves beam scanning, selecting the direction with maximum received power as the AOA. More advanced techniques, such as MUSIC [92, 93] and ESPRIT [94, 95], provide higher angular resolution by exploiting signal subspace properties. A key advantage of angle-based techniques is that they do not require time synchronization between BSs and the device, unlike TOA or TDOA. However, their performance depends heavily on antenna array resolution and the robustness of the estimation algorithm. While higher resolution improves accuracy, it also increases hardware and computational complexity.
II-B3 Fingerprint-based Positioning
@@
Fingerprint-based positioning establishes a mapping between wireless signal features and spatial locations, typically through an offline database or AI model training. A mapping function $\mathcal{F}_{\text{fp}}(\cdot)$ is learned from an offline dataset $D_{\text{offline}}=\{\bm{x}_{i},\bm{p}_{i}\}$, where $\bm{x}_{i}$ denotes fingerprint features and $\bm{p}_{i}$ is the corresponding user location.
As shown in Fig. 4(d), this approach involves two phases: offline and online. In the offline phase, signal characteristics such as RSSI [96, 97], RSRP [98, 99, 100, 101, 102, 103], CSI [104, 105, 106, 107, 108], or TOA [109, 110] are collected at reference points (RPs) to build a fingerprint database or radio map. Features may be extracted from raw data or used directly, and either stored in a database or used to train an AI/ML model that approximates $\mathcal{F}_{\text{fp}}(\cdot)$. In the online phase, real-time measurements $\bm{x}$ from the UE are acquired. Position estimation is then performed by either matching against the fingerprint database or inputting $\bm{x}$ into the trained model, resulting in an estimated position $\hat{\bm{p}}=\mathcal{F}_{\text{fp}}(\bm{x})$.
@@
II-B4 Channel Charting based Positioning
Channel charting is a novel approach leveraging CSI to enable the pseudo-positioning of users by modeling relationships between different channels [111]. It learns a mapping from CSI to a lower-dimensional channel chart, where distances in the latent space represent dissimilarity metrics between channels [112, 113, 114, 115]. These distances indicate user proximity, where greater channel similarity implies closer placement on the channel chart [111]. Besides positioning, channel charting techniques have been explored in various tasks, such as SNR prediction [116], pilot assignment and reuse [117, 118, 119], beam tracking [120], and wireless resource optimization [121]. In contrast to fingerprint-based methods, which depend on pre-collected fingerprints and external RPs to achieve absolute positioning, channel charting aims to uncover latent spatial relationships with minimal reliance on labeled data.
Channel charting, a form of manifold learning, encompasses classical methods such as multidimensional scaling (MDS), Sammon mapping [122], and t-SNE [123], alongside deep learning approaches like Siamese Neural Networks and triplet-based dimensionality reduction. In Siamese-based channel charting systems, dissimilarity metrics define relationships between CSI samples in the latent space. The channel charting learning process aims to ensure that latent space distances preserve the high-dimensional relationships captured by the dissimilarity metrics. A commonly used MDS-based loss function can be given by [113]